1/5 ÷ 4/5 equal to: (a) 4/5 (b) 1/5 (c) 5/4 (d) ¼ 
1/5 ÷ 4/5 equal to: (a) 4/5 (b) 1/5 (c) 5/4 (d) ¼ 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (d) is correct here

= 1/5 ÷ 4/5
= (1/5)/ (4/5)
= (1/5) × (5/4)
= 5/20 ... [divide both numerator and denimonitor by 5]
=1⁄4


Thanks